# Influence of adversarial training on super-resolution turbulence reconstruction

### v4

@@ 4 Effects of adversarial training on in-sample turbulence reconstruction
Both models are tested on in-sample data, statistically matched with the training data, to assess their reconstruction capability. This evaluates the models’ performance under ideal conditions. Figures 2 and 3 compare TKE spectra and the probability density function (PDF) of normalized velocity gradients of the filtered DNS (F-DNS), full-mesh (unfiltered) DNS, TSResNet upsampled fields, and PIESRGAN upsampled fields. The SR models are trained to produce a statistically correct SR field with DNS-like quality. The models are trained and evaluated for the Re90 case using either the box filter (Fig. 2) or spectrally sharp filter (Fig. 3). The primary difference in the training data is the box filter’s artificial attenuation of the near-filter-scale F-DNS fields, which the spectrally sharp filter avoids. The TSResNet recovers the TKE of the resolved scales ($k\gtrsim 32$) but fails to correctly predict the SFS energy distribution for either filter. Conversely, the PIESRGAN is more accurate, deviating from the DNS only for $k\gtrsim 122$ and energy values $E(k)\lesssim 10^{-7}$ for both filters. This behavior is quantitatively confirmed by the PDF of the normalized velocity gradient, where the TSResNet tends to underpredict the large gradients, while the PIESRGAN marginally deviates from the DNS solution. It is worth noting that the TSResNet and the PIESRGAN generator use the same network architecture, while the loss function defers only for the adversarial terms as defined in Sec. 3.2. The PIESRGAN’s more accurate reconstruction is therefore due to its training, rather than the generator’s network depth or loss-function differences. The PIESRGAN architecture improves the root mean squared error (RMSE) by $19\%$ compared to the TSResNet for box-filtered training/testing fields and by $13\%$ for spectrally-sharp filtered training/testing fields.
@@
Figure 3: In-sample (Re90) TKE spectra (left) and PDF of the normalized velocity gradient (right) for spectrally-sharp filtered training/testing fields.
@@
where $u_{i}$ are unfiltered velocity components obtained from either the DNS or the SR fields and $k^{\textsc{sfs}}\equiv(\overline{u_{i}u_{i}}-\overline{u}_{i}\overline{u}_{i})/2$ is the kinetic energy of the subfilter scales [45]. In Eq. 10, the $\overline{\cdot}$ operator indicates filtering (either box or spectrally-sharp filters) and downsampling proportional to the upsampling factor (in the present case, the factor is equal to four). Comparisons are made to the widely used dynamic Smagorinsky model [37, 38],
@@
Similarly, the SFS dissipation rate $\varepsilon^{\textsc{sfs}}=\overline{S}_{ij}\,\tau^{\textsc{sfs}}_{ij}$ is highlighted because it is a crucial property of SFS models [46]. Figure 6 shows the jPDFs of the box-filtered DNS SFS dissipation rate and those of the dynamic-Smagorinsky model (left), the TSResNet-reconstructed field (center), and the PIESRGAN-reconstructed field (right) $\varepsilon^{\textsc{sfs}}$. Here, $\tau^{\textsc{sfs}}_{ij}$ is determined using Eq. 11 for the dynamic Smagorinsky model and Eq. 10 for the SR models and DNS.
@@ 5.1 Different filter kernels for training and testing
Figure 7 compares the TKE spectra and the PDF of normalized velocity gradients for the networks that were trained individually on box-filtered or spectrally-sharp filtered fields exclusively and applied to Gaussian-filtered input data. The exclusively box-filter-trained SR models consistently overestimate the SFS kinetic energy, which results in inaccurate predictions at the high wavenumbers and overpredicted energy at lower wavenumbers than the filter cutoff. Conversely, for models exclusively trained on spectrally sharp filtered data, the opposite behavior is observed when the network is applied to Gaussian-filtered data: both models underpredict the resolved and SFS kinetic energy, with the PIESRGAN model being slightly more accurate than the TSResNet model. Clearly, both SR models fail to extrapolate to filter kernels not seen in training.
@@
One possible approach to address the reduced out-of-sample filter invertibility is to combine an SR architecture with an algebraic approximate deconvolution method [49]. This would first use the ADM to recover the resolved velocity fluctuations, then a PIESRGAN trained for spectrally-sharp filtered fields would extrapolate the subfilter scales. Still, it is important to note that the ADM-deconvoluted field would still differ from an ideal, spectrally-sharp filtered field, especially near the cutoff scale [11, 50]. Thus, the performance of the spectral-trained GAN could suffer from its tendency to invert a known filter. A combination of an SR data-driven model and an algebraic ADM could partially minimize the GAN’s over/underestimation, but it would likely not completely solve the issue.
Another potential solution leverages the GAN’s flexibility to train both in a semi-supervised and unsupervised mode, relying on the underlying “structure” of the data for patterns. This approach will be described subsequently. Both models were trained on a mixture of box-filtered and spectrally-sharp filtered data in an otherwise consistent manner with the training strategy outlined previously (equal number of sub-boxes per filter kernel, randomly selected during the training). This approach is henceforth referred to as multiple-kernel-trained. The discriminator is then trained to distinguish between generated high-resolution fields, upsampled from either box-filtered or spectrally-sharp filtered DNS fields, and the corresponding ground-truth fields. Validating for in-sample F-DNS fields (i.e., obtained with the same filter kernels used for training), the reconstruction performance is comparable to that obtained in Sec. 4. Similar to the previous analysis, the PIESRGAN demonstrates better reconstruction capability than the TSResNet; hence we omit the TKE spectra and PDF of the normalized velocity gradient for brevity. Both models are capable of recognizing and learning both filtering operations that were applied to generate low-resolution training data, suggesting that this ability is not exclusively derived from adversarial training.
@@
Figure 8: 2D slices of instantaneous normalized velocity magnitude and the absolute error of SR fields versus DNS. The normalized absolute error is $\hat{E}=E/\max(E_{\textsc{dns},\textsc{sr}_{1}},E_{\textsc{dns},\textsc{sr}_{2}})$, where $E=\left|u_{\textsc{dns}}-{u}_{\textsc{sr}_{1,2}}\right|$, and the subscripts $1$ and $2$ indicate the multiple-filter-kernel and semi-supervised training approaches, respectively.
While the discriminator does not directly influence the generator’s ability to distinguish between filter kernels, adversarial training can still improve the accuracy of the reconstructed fields. This is an advantage of the GAN-based architecture. Inspired by Bode et al. [19], though differing in purpose and methodology, we now include both labeled (corresponding DNS and F-DNS fields) - as in the previous approach - and unlabeled data (computed LES fields, either with no SFS model or the dynamic Smagorinsky model, as described in Sec. 3) in the training process (referred to LES-training step). Labeled data helps the discriminator differentiate between real and fake samples, while unlabeled data improves the generator’s ability to upsample realistic fields. The discriminator is trained in a supervised manner using only generated and ground-truth fields, and its weights are not updated in the LES-training step, while the generator continues to be trained on randomly sampled LES data. The resulting super-resolved LES fields are then evaluated by the trained discriminator. Since the discriminator is not trained further, the generator loss was reduced to
@@
To isolate the effect the discriminator exerts, we disable either the adversarial or continuity loss in Eq. 12. With this, the generator is trained in a fully unsupervised manner, relying solely on feedback from the continuity equation (1st option) or solely on feedback from the discriminator network (2nd option). Figure 10 compares the possible training options showing joint PDFs of the Gaussian-filtered DNS $\tau^{\textsc{sfs}}_{12}$ (Eq. 10) using only $\mathrm{L_{continuity}}$ loss (left), only $\mathrm{L_{adversarial}}$ loss (center), and both $\mathrm{L_{continuity}}$ and $\mathrm{L_{adversarial}}$ losses (right) during the unsupervised training phase. The PIESRGAN-reconstructed field using only the $\mathrm{L_{adversarial}}$ loss during unsupervised training is statistically similar to that obtained when both losses are employed. Conversely, the reconstructed field obtained using only the $\mathrm{L_{continuity}}$ depicts a bulkier distribution around the diagonal and a predominant shift, resulting in a consistent overestimation of the magnitude of the Reynolds stress. While the reconstruction improvement using only $\mathrm{L_{adversarial}}$ is comparable to that shown in Fig. 9, the training process became more unstable and ultimately resulted in GAN collapse. To address this issue, the batch size and initial learning rate were decreased (the batch size was divided by half and the initial learning rate decreased by an order of magnitude) to improve the training stability [22]. Thus, the continuity loss serves to stabilize and accelerate the training process and does not seem to meaningfully drive reconstruction performance, but further investigations are needed.
@@ 5.2 Different filter sizes for training and testing
Multiple branch employment is promising for image processing [51], though the generator needs prior information on the input size, resulting in the network only being able to handle pre-determined, known input sizes. Small upsampling factors typically result in adequate reconstruction, if incompletely resolved, with larger structures typically accurately captured. Conversely, over-upsampling can lead to significant spurious artifacts [52].
Instead, we use the discriminator to select the appropriate upsampling branch. The GAN is trained for box-filtered fields of different filter sizes ($\overline{\Delta}_{2}$, $\overline{\Delta}_{4}$, and $\overline{\Delta}_{8}$). During training, the correct upsampling branch is enabled based on the known filter size, and unused branches are deactivated. During testing, the generator upsamples the input field along all three branches, producing ×$\times 2$, ×$\times 4$, and ×$\times 8$ fields, which the trained discriminator then uses to calculate an adversarial loss (using Eq. 8 with $\phi_{\textsc{dns}}=0$). The upsampling factor corresponding to the lowest adversarial loss is chosen. Thus, the discriminator learns to recognize high-quality fields during training and can provide highly nonlinear feedback to the generator on the quality of the generated field.
@@
related to the discriminator loss, for $\overline{\Delta}_{2}$, $\overline{\Delta}_{4}$, $\overline{\Delta}_{8}$, and $\overline{\Delta}_{16}$ input data upsampled along the ×$\times 2$, ×$\times 4$, and ×$\times 8$ branches. Higher values indicate higher confidence (by the discriminator) in the quality of the upsampled field. The $\overline{\Delta}_{16}$ input data is out-of-sample. The branch providing the highest quality score is selected as the most-appropriate output for arbitrary input data.
@@
Figure 12 shows normalized velocity-gradient PDFs for PIESRGAN-upsampled fields using the discriminator-selected upsampling branches. It is evident that the discriminator accurately selects the appropriate branch, enabling the modified PIESRGAN to accurately recover both the large scales and the tails of the PDFs. This adaptive upsampling strategy, utilizing the discriminator, is particularly effective in scenarios where the input filter size is not present in the training data. For example, in Tab. 3, the ×$\times 8$ upsampling branch is correctly selected as the most appropriate available branch for the out-of-sample $\overline{\Delta}_{16}$ input data. However, it is important to note that this approach entails higher computational cost compared to using a single upsampling factor (roughly 50 % higher), as more fields are employed.In conclusion, this dynamic upsampling feedback, which can only be provided in the context of a GAN-based architecture, is an effective approach to mitigating the fixed upsampling/filter size limitation.
@@ 6 Extrapolation to higher Reynolds numbers
To test this generalization capability with equal filter kernel and size for both training and testing, thus under ideal conditions, the models trained for the lower $Re_{\lambda}=90$ case (Re90) (Sec. 4) are applied to forced HIT for the higher $Re_{\lambda}=130$ (Re130) and $Re_{\lambda}=350$ (Re350) cases (Tab. 1). Figure 13 compares the TKE spectra for these out-of-sample model applications to the unfiltered and box-filtered ($\overline{\Delta}_{4}$) DNS spectra.
@@ 7 Conclusion
We investigated the influence of adversarial training on super-resolution turbulence reconstruction. GAN-based models (PIESRGANs) are assessed against a standard, supervised CNN-based model (TSResNet). Two novel approaches are introduced that leverage the GAN model’s adversarial training capabilities, enabled by its discriminator network, to enhance the generator network’s accuracy and generalizability to out-of-sample inputs for different filter sizes and filter kernels. The GAN generator and TSResNet have identical model structures, hence their evaluation cost is the same despite the enhanced accuracy of the GAN model.
@@
Adversarial training, while computationally more expensive, is promising to enhance SR model reconstruction of small-scale turbulence. In a priori assessments of the SFS stress tensor, the SR models outperform the widely-used dynamic Smagorinsky model, showing significantly improved statistical alignment with the “true” filtered-DNS SFS stresses. The usefulness of these findings will, however, need to be further verified through a posteriori tests. Of the SR models, the PIESRGAN-reconstructed SFS fields are 64 % better aligned with the filtered-DNS SFS fields than those produced by the TSResNet. This is due entirely to the GAN discriminator’s feedback to the generator; all other training was performed identically. The discriminator’s semi-supervised training enables it to capture more complex, nonlinear relationships between the low- and high-resolution fields than a standard, supervised training approach would permit.
Data-driven super-resolution methods are known to suffer significant performance drops when applied to substantially out-of-sample data. This capability is particularly important to upscale real-world fields. To test this, we investigated two non-consistent applications of the SR models to testing datasets: out-of-sample filter kernels and out-of-sample upscaling ratios. To assess the influence of the training filter kernel, box- and spectrally sharp-filtered SR models were applied to Gaussian-filtered DNS fields. The box/spectral-trained SR models consistently over/underestimated the SFS kinetic energy when applied to Gaussian-filtered fields and produced poorly-aligned jPDFs of $\tau^{\textsc{sfs}}_{12}$. To mitigate this behavior, both architectures were trained for a random collection of fields containing both box- and spectral-filtered data. The ensuing models have slightly greater reconstruction capability than the single-filter-trained models, though both still had a consistent gap to the Gaussian-filtered DNS. The TSResNet was again inaccurate for the subfilter scales, while the PIESRGAN added artifacts at high wavenumbers, with slight underprediction of filter-scale features.

### v5

@@ 4 Effects of adversarial training on in-sample turbulence reconstruction
Both models are tested on in-sample data, statistically matched with the training data, to assess their reconstruction capability. This evaluates the models’ performance under ideal conditions. Figures 2 and 3 compare TKE spectra and the probability density function (PDF) of normalized velocity gradients of the filtered DNS (F-DNS), full-mesh (unfiltered) DNS, TSResNet upsampled fields, and PIESRGAN upsampled fields. The SR models are trained to produce a statistically correct SR field with DNS-like quality. The models are trained and evaluated for the Re90 case using either the box filter (Fig. 2) or spectrally sharp filter (Fig. 3). The primary difference in the training data is the box filter’s artificial attenuation of the near-filter-scale F-DNS fields, which the spectrally sharp filter avoids. The TSResNet recovers the TKE of the resolved scales ($k\gtrsim 32$) but fails to correctly predict the SFS energy distribution for either filter. Conversely, the PIESRGAN is more accurate, deviating from the DNS only for $k\gtrsim 122$ and energy values $E(k)\lesssim 10^{-7}$ for both filters. This behavior is quantitatively confirmed by the PDF of the normalized velocity gradient, where the TSResNet tends to underpredict the large gradients, while the PIESRGAN marginally deviates from the DNS solution. It is worth noting that the TSResNet and the PIESRGAN generator use the same network architecture, while the loss function defers only for the adversarial terms as defined in Sec. 3.2. The PIESRGAN’s more accurate reconstruction is therefore due to its training, rather than the generator’s network depth or loss-function differences. The PIESRGAN architecture improves the root-mean-squared error (RMSE) by $19\%$ compared to the TSResNet for box-filtered training/testing fields and by $13\%$ for spectrally sharp filtered training/testing fields.
@@
Figure 3: In-sample (Re90) TKE spectra (left) and PDF of the normalized velocity gradient (right) for spectrally sharp filtered training/testing fields.
@@
where $u_{i}$ are unfiltered velocity components obtained from either the DNS or the SR fields and $k^{\textsc{sfs}}\equiv(\overline{u_{i}u_{i}}-\overline{u}_{i}\overline{u}_{i})/2$ is the kinetic energy of the subfilter scales [46]. In Eq. (10), the $\overline{\cdot}$ operator indicates filtering (either box or spectrally sharp filters) and downsampling proportional to the upsampling factor (in the present case, the factor is equal to four). Comparisons are made to the widely used dynamic Smagorinsky model [38, 39],
@@
Similarly, the SFS dissipation rate $\varepsilon^{\textsc{sfs}}=\overline{S}_{ij}\,\tau^{\textsc{sfs}}_{ij}$ is highlighted because it is a crucial property of SFS models [47]. Figure 6 shows the jPDFs of the box-filtered DNS SFS dissipation rate and those of the dynamic-Smagorinsky model (left), the TSResNet-reconstructed field (center), and the PIESRGAN-reconstructed field (right) $\varepsilon^{\textsc{sfs}}$. Here, $\tau^{\textsc{sfs}}_{ij}$ is determined using Eq. (11) for the dynamic Smagorinsky model and Eq. (10) for the SR models and DNS.
@@ 5.1 Different filter kernels for training and testing
Figure 7 compares the TKE spectra and the PDF of normalized velocity gradients for the networks that were trained individually on box-filtered or spectrally sharp filtered fields exclusively and applied to Gaussian-filtered input data. The exclusively box-filter-trained SR models consistently overestimate the SFS kinetic energy, which results in inaccurate predictions at the high wave numbers and overpredicted energy at lower wave numbers than the filter cutoff. Conversely, for models exclusively trained on spectrally sharp filtered data, the opposite behavior is observed when the network is applied to Gaussian-filtered data: both models underpredict the resolved and SFS kinetic energy, with the PIESRGAN model being slightly more accurate than the TSResNet model. Clearly, both SR models fail to extrapolate to filter kernels not seen in training.
@@
One possible approach to address the reduced out-of-sample filter invertibility is to combine an SR architecture with an algebraic approximate deconvolution method [50]. This would first use the ADM to recover the resolved velocity fluctuations, then a PIESRGAN trained for spectrally sharp filtered fields would extrapolate the subfilter scales. Still, it is important to note that the ADM-deconvoluted field would still differ from an ideal, spectrally sharp filtered field, especially near the cutoff scale [11, 51]. Thus, the performance of the spectral-trained GAN could suffer from its tendency to invert a known filter. A combination of an SR data-driven model and an algebraic ADM could partially minimize the GAN’s over/underestimation, but it would likely not completely solve the issue.
Another potential solution leverages the GAN’s flexibility to train both in a semisupervised and unsupervised mode, relying on the underlying “structure” of the data for patterns. This approach will be described subsequently. Both models were trained on a mixture of box-filtered and spectrally sharp filtered data in an otherwise consistent manner with the training strategy outlined previously (equal number of subboxes per filter kernel, randomly selected during the training). This approach is henceforth referred to as multiple-kernel-trained. The discriminator is then trained to distinguish between generated high-resolution fields, upsampled from either box-filtered or spectrally sharp filtered DNS fields, and the corresponding ground-truth fields. Validating for in-sample F-DNS fields (i.e., obtained with the same filter kernels used for training), the reconstruction performance is comparable to that obtained in Sec. 4. Similar to the previous analysis, the PIESRGAN demonstrates better reconstruction capability than the TSResNet; hence we omit the TKE spectra and PDF of the normalized velocity gradient for brevity. Both models are capable of recognizing and learning both filtering operations that were applied to generate low-resolution training data, suggesting that this ability is not exclusively derived from adversarial training.
@@
Figure 8: 2D slices of instantaneous normalized velocity magnitude and the absolute error of SR fields versus DNS. The normalized absolute error is $\hat{E}=E/\max(E_{\textsc{dns},\textsc{sr}_{1}},E_{\textsc{dns},\textsc{sr}_{2}})$, where $E=\left|u_{\textsc{dns}}-{u}_{\textsc{sr}_{1,2}}\right|$, and the subscripts $1$ and $2$ indicate the multiple-filter-kernel and semisupervised training approaches, respectively.
While the discriminator does not directly influence the generator’s ability to distinguish between filter kernels, adversarial training can still improve the accuracy of the reconstructed fields. This is an advantage of the GAN-based architecture. Inspired by Bode et al. [19], though differing in purpose and methodology, we now include both labeled (corresponding DNS and F-DNS fields) - as in the previous approach - and unlabeled data (computed LES fields, either with no SFS model or the dynamic Smagorinsky model, as described in Sec. 3) in the training process (referred to as the LES-training step). Labeled data helps the discriminator differentiate between real and fake samples, while unlabeled data improves the generator’s ability to upsample realistic fields. The discriminator is trained in a supervised manner using only generated and ground-truth fields, and its weights are not updated in the LES-training step, while the generator continues to be trained on randomly sampled LES data. The resulting super-resolved LES fields are then evaluated by the trained discriminator. Since the discriminator is not trained further, the generator loss was reduced to
@@
To isolate the effect the discriminator exerts, we disable either the adversarial or continuity loss in Eq. 12. With this, the generator is trained in a fully unsupervised manner, relying solely on feedback from the continuity equation (first option) or solely on feedback from the discriminator network (second option). Figure 10 compares the possible training options showing joint PDFs of the Gaussian-filtered DNS $\tau^{\textsc{sfs}}_{12}$ [Eq. 10] using only $\mathrm{L_{continuity}}$ loss (left), only $\mathrm{L_{adversarial}}$ loss (center), and both $\mathrm{L_{continuity}}$ and $\mathrm{L_{adversarial}}$ losses (right) during the unsupervised training phase. The PIESRGAN-reconstructed field using only the $\mathrm{L_{adversarial}}$ loss during unsupervised training is statistically similar to that obtained when both losses are employed. Conversely, the reconstructed field obtained using only the $\mathrm{L_{continuity}}$ depicts a bulkier distribution around the diagonal and a predominant shift, resulting in a consistent overestimation of the magnitude of the Reynolds stress. While the reconstruction improvement using only $\mathrm{L_{adversarial}}$ is comparable to that shown in Fig. 9, the training process became more unstable and ultimately resulted in GAN collapse. To address this issue, the batch size and initial learning rate were decreased (the batch size was divided by half and the initial learning rate decreased by an order of magnitude) to improve the training stability [22]. Thus, the continuity loss serves to stabilize and accelerate the training process and does not seem to meaningfully drive reconstruction performance, but further investigations are needed.
@@ 5.2 Different filter sizes for training and testing
Multiple branch employment is promising for image processing [52], though the generator needs prior information on the input size, resulting in the network being able to handle only pre determined, known input sizes. Small upsampling factors typically result in adequate reconstruction, if incompletely resolved, with larger structures typically accurately captured. Conversely, over-upsampling can lead to significant spurious artifacts [53].
Instead, we use the discriminator to select the appropriate upsampling branch. The GAN is trained for box-filtered fields of different filter sizes ($\overline{\Delta}_{2}$, $\overline{\Delta}_{4}$, and $\overline{\Delta}_{8}$). During training, the correct upsampling branch is enabled based on the known filter size, and unused branches are deactivated. During testing, the generator upsamples the input field along all three branches, producing ×$\times 2$, ×$\times 4$, and ×$\times 8$ fields, which the trained discriminator then uses to calculate an adversarial loss [using Eq. (8) with $\phi_{\textsc{dns}}=0$]. The upsampling factor corresponding to the lowest adversarial loss is chosen. Thus, the discriminator learns to recognize high-quality fields during training and can provide highly nonlinear feedback to the generator on the quality of the generated field.
@@
related to the discriminator loss, for $\overline{\Delta}_{2}$, $\overline{\Delta}_{4}$, $\overline{\Delta}_{8}$, and $\overline{\Delta}_{16}$ input data upsampled along the ×$\times 2$, ×$\times 4$, and ×$\times 8$ branches. Higher values indicate higher confidence (by the discriminator) in the quality of the upsampled field. The $\overline{\Delta}_{16}$ input data are out-of-sample. The branch providing the highest quality score is selected as the most-appropriate output for arbitrary input data.
@@
Figure 12 shows normalized velocity-gradient PDFs for PIESRGAN-upsampled fields using the discriminator-selected upsampling branches. It is evident that the discriminator accurately selects the appropriate branch, enabling the modified PIESRGAN to accurately recover both the large scales and the tails of the PDFs. This adaptive upsampling strategy, utilizing the discriminator, is particularly effective in scenarios where the input filter size is not present in the training data. For example, in Table 3, the ×$\times 8$ upsampling branch is correctly selected as the most appropriate available branch for the out-of-sample $\overline{\Delta}_{16}$ input data. However, it is important to note that this approach entails higher computational cost compared to using a single upsampling factor (roughly 50 % higher), as more fields are employed.In conclusion, this dynamic upsampling feedback, which can be provided only in the context of a GAN-based architecture, is an effective approach to mitigating the fixed upsampling/filter size limitation.
@@ 6 Extrapolation to higher Reynolds numbers
To test this generalization capability with equal filter kernel and size for both training and testing, thus under ideal conditions, the models trained for the lower $Re_{\lambda}=90$ case (Re90) (Sec. 4) are applied to forced HIT for the higher $Re_{\lambda}=130$ (Re130) and $Re_{\lambda}=350$ (Re350) cases (Table 1). Figure 13 compares the TKE spectra for these out-of-sample model applications to the unfiltered and box-filtered ($\overline{\Delta}_{4}$) DNS spectra.
@@ 7 Conclusion
We investigated the influence of adversarial training on super-resolution turbulence reconstruction. GAN-based models (PIESRGANs) are assessed against a standard, supervised CNN-based model (TSResNet). Two approaches are introduced that leverage the GAN model’s adversarial training capabilities, enabled by its discriminator network, to enhance the generator network’s accuracy and generalizability to out-of-sample inputs for different filter sizes and filter kernels. The GAN generator and TSResNet have identical model structures, hence their evaluation cost is the same despite the enhanced accuracy of the GAN model.
@@
Adversarial training, while computationally more expensive, is promising to enhance SR model reconstruction of small-scale turbulence. In a priori assessments of the SFS stress tensor, the SR models outperform the widely used dynamic Smagorinsky model, showing significantly improved statistical alignment with the “true” filtered-DNS SFS stresses. The usefulness of these findings will, however, need to be further verified through a posteriori tests. Of the SR models, the PIESRGAN-reconstructed SFS fields are 64 % better aligned with the filtered-DNS SFS fields than those produced by the TSResNet. This is due entirely to the GAN discriminator’s feedback to the generator; all other training was performed identically. The discriminator’s semisupervised training enables it to capture more complex, nonlinear relationships between the low- and high-resolution fields than a standard, supervised training approach would permit.
Data-driven super-resolution methods are known to suffer significant performance drops when applied to substantially out-of-sample data. This capability is particularly important to upscale real-world fields. To test this, we investigated two non-consistent applications of the SR models to testing data sets: out-of-sample filter kernels and out-of-sample upscaling ratios. To assess the influence of the training filter kernel, box- and spectrally sharp-filtered SR models were applied to Gaussian-filtered DNS fields. The box/spectral-trained SR models consistently over/underestimated the SFS kinetic energy when applied to Gaussian-filtered fields and produced poorly aligned jPDFs of $\tau^{\textsc{sfs}}_{12}$. To mitigate this behavior, both architectures were trained for a random collection of fields containing both box- and spectral-filtered data. The ensuing models have slightly greater reconstruction capability than the single-filter-trained models, though both still had a consistent gap to the Gaussian-filtered DNS. The TSResNet was again inaccurate for the subfilter scales, while the PIESRGAN added artifacts at high wave numbers, with slight underprediction of filter-scale features.